OpenStudy (anonymous):

WILL FAN AND MEDAL Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of one half. Please just a simple answer

OpenStudy (anonymous):

y − 6 = one half(x + 3) y = one halfx − 6 y + 3 = one half(x − 6) x − 2y = 12

OpenStudy (gabebae):

Point slope form is y - y1 = m ( x - x1 )

OpenStudy (anonymous):

thanks so much

OpenStudy (gabebae):

When you're given two points, x and y values, make sure you change the sign

OpenStudy (gabebae):

Your answer would be the third option

OpenStudy (anonymous):

thank you!